40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13501     Entry time: Wed Jan 3 18:00:46 2018
Author: gautam 
Type: Update 
Category: PonderSqueeze 
Subject: plan of action 

Notes of stuff we discussed @ today's meeting, and afterwards, towards measuring ponderomotive squeezing at the 40m.

  1. Displacement noise requirements
    • Kevin is going to see if we can measure any kind of squeezing on a short timescale by tuning various parameters.
    • Specifically, without requiring crazy ultra low current noise level for the coil driver noise.
  2. Investigate how much actuation range we need for lock acquisition and maintaining lock.
    • Specifically, for DARM.
    • We will measure this by having the arms controlled with ALS in the CARM/DARM basis.
    • Build up a noise budget for this, see how significant the laser noise contribution is.
  3. RC folding mirrors
    • In the present configuration, these are introducing ~2.5% RT loss in the RCs.
    • This affects PRG, and on the output side, measurable squeezing.
    • We want to see if we can relax the requirements on the RC folding mirrors such that we don't have to spend > 20 k$.
    • Specifically, consider spec'ing the folding mirror coatings to only have HR @1064 nm, and take what we get at 532 nm.
    • But still demand tolerances on RoC driven by mode-matching between the RCs and the arm cavities.
  4. ALS with Beat Mouth
    • Use the fiber coupled light from the ends to make the ALS signals.
    • Gautam will update diagram to show the signal chain from end-to-end (i.e. starting at AUX laser, ending at ADC input).
    • Make a noise budget for the same - preliminary analysis suggests a sensing noise floor of ~10 mHz/rtHz.

RXA:

  • For the ALS-DARM budget the idea is that we can do lock acquisition better, so we don't need to care about the acquisition reqs. i.e. we just need to set the ETM coil driver current range based on the DARM in-lock values.
    • To get the coil driver noise to be low enough to detect squeezing we need to use a ~10-15 kOhm series resistor.
    • We assume that all DAC and coil driver input noises can be sufficiently filtered.
    • We are assuming that we don't change the magnet sizes or the number of coil windings in the OSEMs.
    • The noise in the ITMs doesn't matter because we don't use them for any locking activity, so we can easily set the coil driver series resistors to 15 kOhm.
    • We will do the bias for the ETMs and ITMs using some HV circuit (not the existing ones on the coil driver boards) and doing the summation after the main coil driver series resistor. This HV bias module needs to handle the ~ (2 V / 400 Ohm) = 5 mA which is now used. This would require (5 mA) x (15 kOhm) = 60+ V drivers.
  • IF we can get away with doing the ALS beat note with just red (still using GREEN light from the end laser to lock to the arms from the ends), we will not have any requirements for the 532 nm transmission of any optics in the DRMI area.
    • Get some quotes for the new PR/SR mirrors having tight RoC tolerance, high R for 1064, and no spec for 532.
    • Check that the 1-way fiber noise for 1064 nm is < 100 mHz/rHz in the 50-1000 Hz band. If its more, explore putting better acoustic foam around the fiber run.
    • Improve the mode-matching of the IR beam into the fibers at the ends. We want >80% to reduce the noise do to scattering; we don't really care about the amount of light available in the PSL - this is just to reduce the IR-ALS noise.
ELOG V3.1.3-